OpenStudy (anonymous):

Can someone plz help me with this question!! Estimate the area under the curve f(x)= x^2 +2x for 0 less than or eqaul to x less than or equal to 1 by evaluating the riemanna sum with n=8, taking the sample points Ci to be the right-hand end points. Draw a pic illustrate the process

OpenStudy (anonymous):

@IrishBoy123 @zepdrix

OpenStudy (anonymous):



OpenStudy (anonymous):

this is going to take a raft of arithmetic

OpenStudy (anonymous):

divide the interval \((0,1)\) in to 8 parts, each will have length \(\frac{1}{8}=0.125\)

OpenStudy (anonymous):

I have went through my book and notes trying to solve this but do not know how to do this.

OpenStudy (anonymous):

the right hand endpoints will be\[.125,.25,.375,.5,.625,.75,.875,1\]

OpenStudy (anonymous):

or if you perfer \[\frac{1}{8},\frac{2}{8},\frac{3}{8},\frac{4}{8},\frac{5}{8},\frac{6}{8},\frac{7}{8},\frac{8}{8}\]

OpenStudy (anonymous):

you have to evalute the function at each right hand endpoint add them up, and then multiply the result by \[\frac{1}{8}\]

OpenStudy (anonymous):

use a speadsheet or sommat

OpenStudy (anonymous):

it will not be fun, but it is totally doable \[\left(f(0.15)+f(.25)+f(3.75)+f(.5)\\ +f(6.25)+f(.75+f(.875)+f(1)\right)\times .125\]

OpenStudy (anonymous):

damn typo, i am not writing it again

jimthompson5910 (jim_thompson5910):

here's what the graph and rectangles look like (see the attached image)

OpenStudy (anonymous):

that is impressive @jim_thompson5910

jimthompson5910 (jim_thompson5910):

thanks, I had geogebra do all the work

OpenStudy (anonymous):

will it compute the sum too?

OpenStudy (anonymous):

t will not be fun, but it is totally doable (f(0.125)+f(.25)+f(3.75)+f(.5)+f(6.25)+f(.75+f(.875)+f(1))×.125 okay what would f be? I understand the math par.

OpenStudy (anonymous):

and @jim_thompson5910 thanks for the image I will def put that in my notes :)

OpenStudy (anonymous):

\[f(x)=x^2+x\] so for example \[f(.125)=(.125)^2+.125\] lather rinse repeat

OpenStudy (anonymous):

ok that was wrong \[f(x)=x^2+2x\] so \[f(.125)=(.125)^2+2\times .125\]

jimthompson5910 (jim_thompson5910):

yes, what you can do is draw up a spreadsheet like this

OpenStudy (anonymous):

etc etc did i mention it is a pile of arithmetic? and not that useful either since the actual area very easy to find

OpenStudy (anonymous):

looks good take @jim_thompson5910 answer there and multiply it by \(.125\) the length of each interval

jimthompson5910 (jim_thompson5910):

in cell C11, I typed in `Sum[C2:C10]` and for some reason, I went too far, I should have just done `Sum[C2:C9]`, oh well

OpenStudy (anonymous):

yeah u mentioned that it's a lot of arithmetic. But I need to fidn the area this problem is really confusing to me plz explain. and I like the way @jim_thompson5910 did okay I wwill do that @satellite73

OpenStudy (anonymous):

it is the area of each rectangle

OpenStudy (anonymous):

but still not sure how @jim_thompson5910 got those digits there.

OpenStudy (anonymous):

each rectangle has height \(f(x_i)\) for the various \(x_i\) right hand endpoints, and each base is \(.125\) since you divided an interval of length one by eight

OpenStudy (anonymous):

okay x is 1/8, 2/8, 3/8....

OpenStudy (anonymous):